# Image is All You Need: Towards Efficient and Effective Large Language Model-Based Recommender Systems

### v1

@@ Abstract.
Large Language Models (LLMs) have recently emerged as a powerful backbone for recommender systems. Existing LLM-based recommender systems take two different approaches for representing items in natural language, i.e., Attribute-based Representation and Description-based Representation. In this work, we aim to address the trade-off between efficiency and effectiveness that these two approaches encounter, when representing items consumed by users. Based on our interesting observation that there is a significant information overlap between images and descriptions associated with items, we propose a novel method, Image is all you need for LLM-based Recommender system (I-LLMRec). Our main idea is to leverage images as an alternative to lengthy textual descriptions for representing items, aiming at reducing token usage while preserving the rich semantic information of item descriptions. Through extensive experiments, we demonstrate that I-LLMRec outperforms existing methods in both efficiency and effectiveness by leveraging images. Moreover, a further appeal of I-LLMRec is its ability to reduce sensitivity to noise in descriptions, leading to more robust recommendations. Our code is available at https://github.com/rlqja1107/torch-I-LLMRec.
Recommender System, Large Language Models, Sequence modeling
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NY
1. Introduction
@@
Large Language Models (LLMs) have recently demonstrated remarkable performance in various downstream tasks in the Natural Language Processing (NLP) domain, leveraging their strong semantic reasoning and world knowledge. These exceptional capabilities have stimulated researchers to explore their potential and broaden the application to diverse domains (Kim et al., 2024b; Chen et al., 2024; Kim et al., 2024a; Xie et al., 2024). In recommender systems, substantial efforts have also been made to utilize their rich semantic knowledge. Especially, recent studies explore the replacement of traditional collaborative filtering models (e.g., SASRec (Kang and McAuley, 2018) and LightGCN (He et al., 2020)) with LLMs as the backbone for recommendations (Lin et al., 2024; Bao et al., 2023; Zheng et al., 2024; Tan et al., 2024). To this end, they typically transform each item in a user’s interaction history from a traditional numerical ID into natural language (e.g., titles) and arrange them into sequences within the input prompt.
Figure 1. (a) Histogram of the input token length required to represent a user’s item interaction history across different item expression approaches. (b) Recommendation performance (Hit@5) and Inference Time (seconds/100 users) for different item representation approaches. We use the Amazon Sports dataset for analysis.
@@
The key to LLM-based recommender systems lies in effectively representing items in natural language to capture user preferences based on LLM’s comprehensive understanding of user interaction history. In this regard, existing studies that represent items in natural language can be categorized into the following two main approaches: 1) Attribute-Based Representation approach is to combine simple attributes such as brand and category with the item title to represent the item (Bao et al., 2023; Li et al., 2023b; Tan et al., 2024; Lin et al., 2024). For example, TALLRec (Bao et al., 2023) represents items with their title for user preference modeling, while RECFORMER (Li et al., 2023b) enhances user preference understanding by adding brand and category attributes with attribute type-specific tokens in the prompt. Similarly, TransRec (Lin et al., 2024) incorporates attribute information alongside the title in the prompt to facilitate a multifaceted understanding of items. 2) Description-Based Representation approach is to use detailed item descriptions to preserve rich textual item semantics so that the LLM could capture user preference in a more fine-grained manner. For example, TRSR (Zheng et al., 2024) summarizes full descriptions using a large-scale LLM and provides them to a small-scale LLM for recommendation, thereby overcoming the limited item semantics provided by attributes only and addressing the input length constraints of LLMs (e.g., 1,024 tokens for GPT-2 (Radford et al., 2019)).
@@
However, we argue that the two aforementioned approaches inherently face a trade-off between efficiency and effectiveness when representing items. Specifically, the Attribute-Based Representation approach is efficient since it requires fewer tokens to add item attributes than to add the entire item descriptions, which shortens the input length for the LLM [^1] Since LLMs are built on transformers, their computational complexity grows quadratically as the input token length increases.. However, it sacrifices the effectiveness of understanding user preferences due to the limited item semantics that can be contained in relatively short attributes. On the other hand, the Description-Based Representation approach is less efficient due to the longer input length required for detailed item descriptions, whereas the effectiveness of understanding user preferences is improved by utilizing richer textual semantics of items. In fact, as shown in Figure 1(a), the input length required for LLMs to represent a user’s item interaction history significantly varies depending on how the item is represented. That is, the Description-Based Representation approach (i.e., Summarized description [^2] To summarize the full description, we adopt an approach similar to TRSR (Zheng et al., 2024) and use a similar prompt. For more details, please refer to Appendix B. and Full description) demands much longer input token lengths than the Attribute-Based Representation approach (i.e., Attribute), resulting in the increased computational complexity.
To better understand this trade-off, in Figure 1(b), we extend our analysis beyond a simple comparison of input token length and examine the computational complexity of different item representation approaches (i.e., efficiency) along with their recommendation performance (i.e., effectiveness). Note that we use the same LLM backbone (i.e., Sheared LLaMA-2.7B (Xia et al., 2023)) and adopt the same recommendation protocol for fair comparisons (See Appendix C regarding the details of recommendation protocol). We observe that the Attribute-based Representation approach (blue circle) achieves the inference time more than 2.5 times faster than the Description-based Representation approach (green and black). However, this comes at the expense of a performance reduction of over 13% due to the lack of rich item semantics contained in attributes compared with descriptions. Furthermore, summarizing full descriptions reduces inference time but compromises performance due to the information loss during the summarization process. In conclusion, the trade-off can be summarized as: richer (i.e., longer) item representation provided to the LLM improves effectiveness but decreases efficiency. However, this trade-off is unavoidable by nature if items are to be represented with natural language.
@@
Figure 2. Cosine similarity between item image-description pairs in Amazon Sport and Art datasets and image-caption pairs in the COCO dataset using CLIP (Radford et al., 2021).
@@
In this paper, we focus on addressing this trade-off based on our interesting observation that there is a significant information overlap between images and descriptions associated with items. Specifically, when we measure the similarity between item image-description pairs in a real-world e-commerce dataset (i.e., Amazon Sport and Art (Ni et al., 2019)) using a vision-language model (i.e., CLIP (Radford et al., 2021)) in which image and language spaces are jointly embedded, we found that the similarity is surprisingly high. Specifically, Figure 2 shows that the average similarity for the Amazon Sport and Art datasets is around 0.31. To better understand this similarity level, we conducted the same experiment with the COCO caption dataset (Lin et al., 2014), a well-curated image-caption pair dataset widely used in vision-language research. Specifically, we measured similarities for both positive (i.e., image-caption) and negative (i.e., image-randomly sampled unrelated caption) pairs. Given that the negative pairs have an average similarity of only 0.07, the positive pairs’ average similarity of around 0.26 can be considered an indicator of high similarity. Furthermore, the similarity value of 0.31 between item image-description pairs, which is even higher, highlights a significant information overlap between images and their associated descriptions [^3] Such an observation was consistently observed across other real-world datasets as well as another CLIP variant model (See Appendix D)..
@@
Building on this observation, we propose a novel method, Image is all you need for LLM-based Recommender system (I-LLMRec), which addresses both efficiency and effectiveness of existing LLM-based recommender systems by leveraging images as an alternative to lengthy textual descriptions for the item expression, aiming at reducing token usage while preserving the rich semantic information of item descriptions [^4] Some text-specific information that is missing in an image may be lost, but we find that in Section 4.2, this information has surprisingly little impact on recommendations.. The main technical challenge lies in the misalignment between the item image space and the language space. To this end, we adopt a learnable adaptor for visual features followed by a technique to bridge the gap between the two spaces (i.e., Image-LLM Alignment mode (ILA)), which facilitates the training of the adaptor with carefully crafted prompts tailored to the recommendation context. This ensures the LLM to capture rich item semantics through images with only a few tokens, thereby effectively and efficiently capturing user preferences.
@@
Through extensive experiments on various real-world datasets, we demonstrate that our proposed method using images instead of item descriptions is superior in both effectiveness and efficiency. Specifically, I-LLMRec improves inference speed by approximately 2.93 times compared to the Description-based Representation approach, while achieving a 22% performance improvement over the Attribute-based Representation approach (Figure 1(b)). As a further appeal, I-LLMRec facilitates robust recommendations by mitigating sensitivity to noise associated with item descriptions as it relies on images. Our main contributions can be summarized as follows:
@@ 2. Preliminary
In this section, we first introduce the task formulation and then examine the complexity of different item expression approaches.
@@
Task Formulation. Throughout this paper, we mainly focus on sequential recommendation, as it closely aligns with real-world scenarios (Tian et al., 2022; Xie et al., 2022). Let $\mathcal{U}$ and $\mathcal{I}$ denote the set of users and items, respectively. A user $u\in\mathcal{U}$ has the historical item interaction sequence denoted as $\mathcal{S}_{u}$=[$i_{1},i_{2},...,i_{k},...,i_{|\mathcal{S}_{u}|}$], where $i_{k}$ denotes the $k$-th interacted item and $|\mathcal{S}_{u}|$ is the number of items in the interaction sequence. The goal of this task is, for each user $u$, to predict the next item $i_{|\mathcal{S}_{u}|+1}$ to be consumed by the user based on the user interaction history $\mathcal{S}_{u}$.
@@
Representing Item Semantics. With the advancement of LLMs for recommender systems, there is a growing need to help them understand the semantics of items to capture user preferences. In this regard, we categorize the multiple types of information to recommend an item $i$’s semantics into [$\mathbf{I}_{i}$, $\mathbf{D}_{i}$, $\mathbf{A}_{i}$], where $\mathbf{I}_{i}$, $\mathbf{D}_{i}$, and $\mathbf{A}_{i}$ are an item image, textual description [^5] To avoid confusion, we will henceforth refer to the description as summarized description instead of full description unless stated otherwise., and attribute (e.g., brand and category), respectively.
@@
Comparison of complexity. Let $f(\cdot)$ denote the transformation function for the LLM to convert the representation of item semantics into input tokens, and let $|f(\cdot)|$ denote the number of input tokens. Specifically, for $\mathbf{D}_{i}$ and $\mathbf{A}_{i}$ that are in natural language, $f$ refers to tokenizing the natural language and converting them into the word embeddings. Meanwhile, for an image $\mathbf{I}_{i}$, $f$ involves extracting the visual features using a pretrained vision encoder (e.g., CLIP-ViT (Radford et al., 2021)), followed by an adaptor to resize their feature dimensions for compatibility with the LLM (Liu et al., 2024b; Li et al., 2023a). The complexity of LLM operations for each representation is $O((|f(\mathbf{I}_{i})||\mathcal{S}_{u}|)^{2}d)$, $O((|f(\mathbf{D}_{i})||\mathcal{S}_{u}|)^{2}d)$, and $O((|f(\mathbf{A}_{i})||\mathcal{S}_{u}|)^{2}d)$, where $|f(\mathbf{D}_{i})|>>|f(\mathbf{A}_{i})|$. This is derived based on the complexity of Transformer, i.e., $O(n^{2}d)$, where $n$ is the input token length and $d$ is the feature dimension of the LLM. It reveals that the complexity of the description-based approach is not merely being incurred with a single item; rather, it increases with the user sequence length $|\mathcal{S}_{u}|$, and its complexity even grows quadratically, making this approach impractical. Therefore, in this work, we propose to lower the complexity of expressing an item by leveraging its associated image using only a single token (i.e., $|f(\mathbf{D}_{i})|\gg|f(\mathbf{A}_{i})|>|f(\mathbf{I}_{i})|=1$ ) [^6] In average, $|f(\mathbf{D}_{i})|$ and $|f(\mathbf{A}_{i})|$ are approximately 160 and 10 tokens, respectively. while preserving the rich semantics of the item description. This is based on our observation that there is a significant information overlap between the item image-description pairs (Figure 2).
3. Proposed Method: I-LLMRec
@@
In this section, we provide a detailed explanation of I-LLMRec. After introducing the initial recommendation-related task prompt, we begin by encoding the images of items consumed by users, and mapping them to an LLM through a learnable adaptor (Section 3.1). However, since these mapped visual features are not inherently aligned with the language space, we propose the Image-LLM Alignment (ILA) module that trains the adaptor with carefully crafted prompts tailored for the recommendation context, resulting in effectively bridging the alignment gap (Section 3.2). Meanwhile, we propose the Image-based REtrieval (IRE) module, a training approach that enables the LLM to directly perform recommendations from the item corpus based on the visual features (Section 3.3). Finally, we describe the overall training objectives and inference for recommendation (Section 3.4 and 3.5). The overall framework of I-LLMRec is shown in Figure 3.
3.1. Mapping Item Images to LLM
@@
Recall that, as shown in Figure 2, the item image contains rich item semantics of the lengthy description. Therefore, in this section, our goal is to map items consumed by users to an LLM using only a few tokens, allowing us to capture user preference efficiently and effectively. Specifically, given the interaction history of user $u$, i.e., $[i_{i},i_{2},...,i_{|\mathcal{S}_{u}|-1}]$, and a target item $i_{|\mathcal{S}_{u}|}$, we convert the interaction history into a sequence of item images $[\mathbf{I}_{i_{1}},\mathbf{I}_{i_{2}},...,\mathbf{I}_{i_{|\mathcal{S}_{u}|-1}}]\in\mathbb{R}^{(|\mathcal{S}_{u}|-1)\times H\times W\times 3}$, where $H$ and $W$ are the image height and width, respectively. Then, we adopt a pretrained visual encoder $V$ of a vision-language model to extract the visual features of each item, defined as $v_{i}=V(\mathbf{I}_{i})\in\mathbb{R}^{d_{v}}$. This process makes the sequence of visual features $[v_{i_{1}},v_{i_{2}},...,v_{i_{|\mathcal{S}_{u}|-1}}]\in\mathbb{R}^{(|\mathcal{S}_{u}|-1)\times d_{v}}$, where $d_{v}$ is the dimension of the visual feature.
3.1.1. Adaptor network for visual features.
@@
However, since the visual feature dimension $d_{v}$ is different from that of the LLM feature $d$, and their spaces are inherently misaligned, we introduce an adaptor network $M:\mathbb{R}^{d_{v}}\rightarrow\mathbb{R}^{d}$ to make visual features compatible with the LLM’s feature dimension, allowing $v_{i}$ to be interactive with word embeddings within the LLM’s input layer. More precisely, we process each item’s visual feature $v_{i}$ through the adaptor $M$, transforming it into a sequence of visual features that match the dimensionality of the LLM. Formally, we the sequence of mapped visual features is defined as:
(1) $\mathcal{\bar{S}}_{u}^{\mathbf{I}}=[\bar{v}_{i_{1}},\bar{v}_{i_{2}},...,\bar{v}_{i_{|\mathcal{S}_{u}|-1}}],\bar{v}_{i}=M(v_{i})\in\mathbb{R}^{d}.$

### v2

# arXiv:2503.06238v2[cs.IR] 28 Feb 2026

@@ Abstract
Large Language Models (LLMs) have recently emerged as a powerful backbone for recommender systems. Existing LLM-based recommender systems take two different approaches for representing items in natural language, i.e., Attribute-based Representation and Description-based Representation. In this work, we aim to address the trade-off between efficiency and effectiveness that these two approaches encounter, when representing items consumed by users. Based on our observation that there is a significant information overlap between images and descriptions associated with items, we propose a novel method, Image representation for LLM-based Recommender system (I-LLMRec). Our main idea is to leverage images as an alternative to lengthy textual descriptions for representing items, aiming at reducing token usage while preserving the rich semantic information of item descriptions. Through extensive experiments on real-world Amazon datasets, we demonstrate that I-LLMRec outperforms existing methods that leverage textual descriptions for representing items in both efficiency and effectiveness by leveraging images. Moreover, a further appeal of I-LLMRec is its ability to reduce sensitivity to noise in descriptions, leading to more robust recommendations. Our code is available at https://github.com/rlqja1107/torch-I-LLMRec.
1 Introduction
@@
LLMs, which have recently shown remarkable performance in various NLP tasks by leveraging strong semantic reasoning and world knowledge, have inspired research into their application across diverse domains, including recommender systems  (Kim et al., 2024a; Chen et al., 2024; Kim et al., 2024b; Xie et al., 2024). Especially, recent studies explore the replacement of traditional collaborative filtering models (e.g., SASRec (Kang and McAuley, 2018)) with LLMs as the backbone for recommendations  (Lin et al., 2024b; Bao et al., 2023; Tan et al., 2024). To this end, they typically transform each item in a user’s interaction history from a traditional numerical ID into natural language (e.g., titles) and arrange them into sequences within the input prompt.
@@
The key to LLM-based recommender systems lies in effectively representing items in natural language to capture user preferences based on LLM’s comprehensive understanding of user interaction history. In this regard, existing studies that represent items in natural language can be categorized into the following two main approaches: 1) Attribute-Based Representation approach is to combine simple attributes such as brand and category with the item title to represent the item (Bao et al., 2023; Li et al., 2023a; Tan et al., 2024). For example, TALLRec (Bao et al., 2023) and TransRec (Lin et al., 2024b) enhance the understanding of user preferences by adding attributes, thereby facilitating a multifaceted understanding of items 2) Description-Based Representation approach is to use detailed item descriptions [^1] While attributes are the high-level, general features in a few keywords (e.g., Apple), descriptions generally provide item-specific details (e.g., It is a slim metallic body, 13-inch Liquid Retina, and black keyboard…). to preserve rich textual item semantics so that the LLM could capture user preference in a more fine-grained manner. For example, TRSR (Zheng et al., 2024b) summarizes full descriptions and provides them to the LLM for recommendation, thereby overcoming the limited item semantics provided by attributes only and addressing the input length constraints of LLMs. Overall, these approaches demonstrate the potential of natural language in enriching item representations for LLM-based recommendation.
@@
However, we argue that when representing items in natural language, there is an inherent trade-off between efficiency and effectiveness. Specifically, the Attribute-Based Representation approach is efficient since it requires fewer tokens to add item attributes than to add the entire item descriptions, which shortens the input length for the LLM. However, it sacrifices the effectiveness of understanding user preferences due to the limited item semantics that can be contained in relatively short attributes. On the other hand, the Description-Based Representation approach is less efficient due to the longer input length required for detailed item descriptions, whereas the effectiveness of understanding user preferences is improved by incorporating richer textual semantics of items. In fact, as shown in Figure 1(a), the input length required for LLMs to represent a user’s item interaction history significantly varies depending on how the item is represented. That is, the Description-Based Representation approach (i.e., Summarized description [^2] To summarize the full description, we adopt an approach similar to TRSR (Zheng et al., 2024b) and use a similar prompt. For more details, please refer to Appendix A. and Full description) demands much longer input token lengths than the Attribute-Based Representation approach (i.e., Attribute), resulting in the increased computational complexity.
@@
Figure 1: (a) Histogram of the input token length required to represent a user’s item interaction history across different item expression approaches. (b) Recommendation performance (Hit@5) and Inference Time (seconds/100 users) for different item representation approaches. For (a) and (b), we use the Amazon Sports dataset for analysis. (c) Cosine similarity between item image-description pairs in Amazon Sport and Art datasets and image-caption pairs in the COCO dataset using CLIP.
To better understand this trade-off, in Figure 1(b), we extend our analysis beyond a simple comparison of input token length and examine the computational complexity of different item representation approaches (i.e., efficiency) along with their recommendation performance (i.e., effectiveness). Note that we use the same LLM backbone (i.e., Sheared LLaMA-2.7B (Xia et al., 2023)) and adopt the same recommendation protocol for fair comparisons (See Appendix B regarding the details of the recommendation protocol). We observe that the Attribute-based Representation approach (blue circle) achieves the inference time more than 2.5 times faster than the Description-based Representation approach (green and black). However, this comes at the expense of a performance reduction of over 13% due to the lack of rich item semantics contained in attributes compared with descriptions. Furthermore, summarizing full descriptions reduces inference time but compromises performance due to the information loss during the summarization. In conclusion, the trade-off can be summarized as: richer item representation provided to the LLM improves effectiveness but decreases efficiency. However, this trade-off is unavoidable by nature if items are represented with natural language.
@@
In this paper, we focus on addressing this trade-off based on our observation that there is a significant information overlap between images and descriptions associated with items. Specifically, when we measure the similarity between item image-description pairs in a real-world e-commerce dataset (i.e., Amazon Sport and Art (Ni et al., 2019)) using a vision-language model (i.e., CLIP (Radford et al., 2021)) in which image and language spaces are jointly embedded, we found that the similarity is surprisingly high. Specifically, Figure 1(c) shows that the average similarity for the Amazon Sport and Art datasets is around 0.31. To better understand this similarity level, we conducted the same experiment with the COCO caption dataset (Lin et al., 2014), a well-curated image-caption pair dataset widely used in vision-language research. Specifically, we measured similarities for both positive (i.e., image-caption) and negative (i.e., image-randomly sampled unrelated caption) pairs. Given that the negative pairs have an average similarity of only 0.07, the positive pairs’ average similarity of around 0.26 can be considered an indicator of high similarity. Furthermore, the similarity value of 0.31 between item image-description pairs, which is even higher, highlights a significant information overlap between images and their associated descriptions [^3] Such an observation was consistently observed across other real-world datasets as well as another CLIP variant model (See Appendix C and E.4)..
@@
Building on this observation, we propose Image representation for LLM-based Recommender system (I-LLMRec), which addresses both efficiency and effectiveness of existing LLM-based recommender systems by leveraging images as an alternative to lengthy textual descriptions for the item representation, aiming at reducing token usage while preserving the rich semantic information of item descriptions [^4] Some text-specific information that is missing in an image may be lost, but we find that in Section 4.2, this information has surprisingly little impact on recommendations.. The main technical challenge lies in the misalignment between the item image space and the language space. To this end, we adopt a learnable adaptor for visual features followed by a technique to bridge the gap between the two spaces (i.e., Recommendation-oriented Image-LLM Semantic Alignment (RISA) module), which facilitates the training of the adaptor with carefully crafted prompts tailored to the recommendation context. This ensures the LLM to capture rich item semantics through images with only a few tokens, thereby effectively and efficiently capturing user preferences.
@@
Through extensive experiments on real-world Amazon datasets, we demonstrate that our proposed method using images instead of item descriptions is superior in both effectiveness and efficiency. Specifically, I-LLMRec improves inference speed by approximately 2.93 times compared to the Description-based Representation approach, while achieving a 22% performance improvement over the Attribute-based Representation approach (Figure 1(b)). As a further appeal, I-LLMRec facilitates robust recommendations by mitigating sensitivity to noise associated with item descriptions as it leverages images. Our main contributions can be summarized as follows:
@@ 2 Preliminaries
Here, we introduce the task formulation and examine the complexity of different item expression approaches.
@@
Task Formulation. Throughout this paper, we mainly focus on sequential recommendation, as it closely aligns with real-world scenarios (Tian et al., 2022; Xie et al., 2022). Let $\mathcal{U}$ and $\mathcal{I}$ denote the set of users and items, respectively. A user $u\in\mathcal{U}$ has the historical item interaction sequence denoted as $\mathcal{S}_{u}$$i_{1},i_{2},...,i_{k},...,i_{|\mathcal{S}_{u}|}$], where $i_{k}$ denotes the $k$-th interacted item and $|\mathcal{S}_{u}|$ is the number of items in the interaction sequence. The goal of this task is, for each user $u$, to predict the next item $i_{|\mathcal{S}_{u}|+1}$ to be consumed by the user based on the user interaction history $\mathcal{S}_{u}$.
@@
Representing Item Semantics. With the advancement of LLMs for recommender systems, there is a growing need to help them understand the semantics of items to capture user preferences. In this regard, we categorize the multiple types of information to recommend an item $i$’s semantics into $\mathbf{I}_{i}$, $\mathbf{D}_{i}$, $\mathbf{A}_{i}$], where $\mathbf{I}_{i}$, $\mathbf{D}_{i}$, and $\mathbf{A}_{i}$ are an item image, textual description [^5] To avoid confusion, we will henceforth refer to the description as summarized description instead of full description unless stated otherwise., and attribute, respectively.
@@
Comparison of Complexity. Let $f(\cdot)$ denote the transformation function for the LLM to convert the representation of item semantics into input tokens, and let $|f(\cdot)|$ denote the number of input tokens. Specifically, for $\mathbf{D}_{i}$ and $\mathbf{A}_{i}$ that are in natural language, $f$ refers to tokenizing the natural language and converting them into the word embeddings. Meanwhile, for an image $\mathbf{I}_{i}$, $f$ involves extracting the visual features using a pretrained vision encoder (e.g., CLIP-ViT (Radford et al., 2021)), followed by an adaptor to resize their feature dimensions for compatibility with the LLM (Liu et al., 2024a; Li et al., 2023b). The complexity of LLM operations for each representation is $O((|f(\mathbf{I}_{i})||\mathcal{S}_{u}|)^{2}d)$, $O((|f(\mathbf{D}_{i})||\mathcal{S}_{u}|)^{2}d)$, and $O((|f(\mathbf{A}_{i})||\mathcal{S}_{u}|)^{2}d)$, where $|f(\mathbf{D}_{i})|>>|f(\mathbf{A}_{i})|$. This is derived based on the complexity of Transformer, i.e., $O(n^{2}d)$, where $n$ is the input token length and $d$ is the feature dimension of the LLM. It reveals that the complexity of the description-based approach is not merely being incurred with a single item; rather, it increases with the user sequence length $|\mathcal{S}_{u}|$, and its complexity even grows quadratically, making this approach impractical. Therefore, in this work, we propose to lower the complexity of expressing an item by leveraging its associated image using only a single token (i.e., $|f(\mathbf{D}_{i})|\gg|f(\mathbf{A}_{i})|>|f(\mathbf{I}_{i})|=1$ ) [^6] In average, $|f(\mathbf{D}_{i})|$ and $|f(\mathbf{A}_{i})|$ are approximately 160 and 10 tokens, respectively. while preserving the rich semantics of the item description. This is based on our observation that there is a significant information overlap between the item image-description pairs (Figure 1(c)).
3 Proposed Method: I-LLMRec
@@
In this section, we provide a detailed explanation of I-LLMRec. After introducing the initial recommendation-related task prompt, we begin by encoding the images of items consumed by users, and mapping them to an LLM through a learnable adaptor (Section 3.1). However, since these mapped visual features are not inherently aligned with the language space, we propose the Recommendation-oriented Image-LLM Semantic Alignment (RISA) module that trains the adaptor with carefully crafted prompts tailored for the recommendation context, resulting in effectively bridging the alignment gap (Section 3.2). Meanwhile, we propose the REtrieval-based Recommendation via Image features (RERI) module, a training approach that enables the LLM to directly perform recommendations from the item corpus based on the visual features (Section 3.3). Finally, we describe the overall training objectives and inference for recommendation (Section 3.4). Figure 2 shows the overall framework of I-LLMRec.
3.1 Mapping Item Images to an LLM
@@
In this section, our goal is to map items consumed by users to an LLM using only a few tokens, allowing us to capture user preference efficiently and effectively. Specifically, given the interaction history of user $u$, i.e., $i_{1},i_{2},...,i_{|\mathcal{S}_{u}|-1}]$, and a target item $i_{|\mathcal{S}_{u}|}$, we convert the interaction history into a sequence of item images $\mathbf{I}_{i_{1}},\mathbf{I}_{i_{2}},...,\mathbf{I}_{i_{|\mathcal{S}_{u}|-1}}]\in\mathbb{R}^{(|\mathcal{S}_{u}|-1)\times H\times W\times 3}$, where $H$ and $W$ are the image height and width, respectively. Then, we adopt a pretrained visual encoder $V$ of a vision-language model to extract the visual features of each item, defined as $v_{i}=V(\mathbf{I}_{i})\in\mathbb{R}^{d_{v}}$. This process makes the sequence of visual features $v_{i_{1}},v_{i_{2}},...,v_{i_{|\mathcal{S}_{u}|-1}}]\in\mathbb{R}^{(|\mathcal{S}_{u}|-1)\times d_{v}}$, where $d_{v}$ is the dimension of the visual feature.
@@
Adaptor network for visual features. However, since the visual feature dimension $d_{v}$ is different from that of the LLM feature $d$, and their spaces are inherently misaligned, we introduce an adaptor network $M:\mathbb{R}^{d_{v}}\rightarrow\mathbb{R}^{d}$ to make visual features compatible with the LLM’s feature dimension, allowing $v_{i}$ to be interactive with word embeddings within the LLM’s input layer. More precisely, we process each item’s visual feature $v_{i}$ through the adaptor $M$, transforming it into a sequence of visual features that match the dimensionality of the LLM. Formally, the sequence of mapped visual features is defined as $\mathcal{\bar{S}}_{u}^{\mathbf{I}}=[\bar{v}_{i_{1}},\bar{v}_{i_{2}},...,\bar{v}_{i_{|\mathcal{S}_{u}|-1}}]$, where $\bar{v}_{i}=M(v_{i})\in\mathbb{R}^{d}.$